Word Problem
Wild and Crazy Productions wants to purchase a new CD producing machine. Its cost is $5000.00. If their labor and other unit costs are $5.00 per CD and they can sell the CD for $10.00 each, how many must they manufacture before they pay for their new machine?

Found 2 solutions by stanbon, checkley75:
Answer by stanbon(75887)

The profit on each CD is $5.00
5000/5 = 1000 CD's
Selling 1000 CD's will give them back their $5000.
Cheers,
Stan H.

Answer by checkley75(3666)

5000=(10-5)x
5000=5x
x=5000/5
x=1000 CD's are required to pay for the machine.
proof:
1000*10-1000*5=5000
10000-5000=5000
5000=5000